The length of a rectangular shaped box is 12 cm more than the height, and the width is 3 cm less than the height. The volume is 324 cm^3. Find the length, width, and height.
Answer by mananth(16946)

The length of a rectangular shaped box is 12 cm more than the height, and the width is 3 cm less than the height. The volume is 324 cm^3. Find the length, width, and height.
...
h= x cm
L=x+12
width = x-3
...
Volume V = 324 cm^3
...
Volume V = L*W*H
x(x+12)(x-3)=324
x(x^2+9x-36)=324
x^3+9x^2-36x-324=0
(x-6)(x+6)(x+9)=0
take the positive root
h = 6
L= 18
W=3
...
m.ananth@hotmail.ca